OpenStudy (anonymous):

Find the 14th term in the sequence for 32, 16, 8, 4, 2, 1, 1/2, 1/4

OpenStudy (anonymous):

Do you know which formula to apply?

OpenStudy (anonymous):

14th term = first term * (ratio)^14 = 32 * (0.5)^14

OpenStudy (anonymous):

the formula is An=A1r ^n-1

OpenStudy (anonymous):

Yep, now plug the numbers in :)

OpenStudy (anonymous):

yes, i did that, no how do i find the 14th term/ answer

OpenStudy (anonymous):

a_1 = 32 r = ?

OpenStudy (anonymous):

r=2

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

Oops, 32r = 16 -> r = 16/32 = ?

OpenStudy (anonymous):

whats the 14th term

OpenStudy (anonymous):

You can't find a_n, if you don't have r yet!

OpenStudy (anonymous):

R is 2.. R is the difference in the sequence

OpenStudy (anonymous):

It's incorrect, because the sequence decreasing! So you can't say r = 2 ( double increasing)

OpenStudy (anonymous):

Its Geometric you divide 2 every time so its 2

OpenStudy (anonymous):

divide by 2 => r = 1/2, not 2!

OpenStudy (anonymous):

okay so what the 14th sequence

OpenStudy (anonymous):

Now, plug a_1 and r into the forumula!

OpenStudy (anonymous):

i know! what is the answer

OpenStudy (anonymous):

Do you have YOUR answer, I'll check it!

OpenStudy (anonymous):

nevermind.... i dont know it, so ill just get an F.

OpenStudy (anonymous):

Why o.O all you do is plug 32* (.5)^13

OpenStudy (anonymous):

What's so hard to plug the number into your calculator???